Exercise 9.4 Depreciation Methods (LO9-3)

On January 2, 2018, Jatson Corporation acquired a new machine with an estimated useful life of five years. The cost of the equipment was $50,000 with an estimated residual value of $4,000.

a-1. Prepare a complete depreciation table under the straight-line method. Assume that a full year of depreciation was taken in 2018.

a-2. Prepare a complete depreciation table under the 200 percent declining-balance method. Assume that a full year of depreciation was taken in 2018.

a-3. Prepare a complete depreciation table under the 150 percent declining-balance with a switch to straight-line when it will maximize depreciation expense. Assume that a full year of depreciation was taken in 2018.
Answer #1

A-1) Straight line method of depreciation :-

Depreciation per year = (50,000 - 4,000)/5 = $9,200

Year Opening balance Depreciation Closing balance
2018 50,000 9,200 40,800
2019 40,800 9,200 31,600
2020 31,600 9,200 22,400
2021 22,400 9,200 13,200
2022 13,200 9,200 4,000

A-2) 200 percent declining balance method :-

Depreciation rate = (1/5) x 200% = 40%

Year Opening balance Depreciation rate Depreciation expense Closing balance
2018 50,000 40% 20,000 30,000
2019 30,000 40% 12,000 18,000
2020 18,000 40% 7,200 10,800
2021 10,800 40% 4,320 6,480
2022 6,480 38.27 2,480 4,000

2022 year depreciation lower of (6,480 x 40%) or (6,480 - 4,000)

= 2,592 or 2,480

Therefore, 2,480 is Depreciation expense, rate = 2,480/6,480 = 38.27%

A-3) 150 percent declining balance method :-

Depreciation rate = (1/5) x 150% = 30%

Year Opening balance Depreciation expense Closing balance
2018 50,000 15,000 35,000
2019 35,000 10,500 24,500
2020 24,500 7,350 17,150
2021 17,150 6,575 10,575
2022 10,575 6,575 4,000

Depreciation expense from year 2021 can be maximised with straight line method.

Opening Balance in 2021 = 17,150

Straight line depreciation = (17,150 - 4,000)/2 = 6,575

150% depreciation = 17,150 x 30% = 5,145

Therefore straight line method gives higher depreciation.